OpenStudy (anonymous):

find all solutions to tan(theta)=(square root)3

OpenStudy (anonymous):

\[find all solutions \to \tan \theta=\sqrt{3}\]

OpenStudy (anonymous):

\[tan(\theta)=\sqrt{3}\] \[\frac{1}{2} \div\sqrt{3}/2=\sqrt{3}\] \[tan(\theta)=sin(\theta)/cos(\theta)\]

OpenStudy (anonymous):

sin(30)=1/2 cos(30)= sqrt[3]/2

OpenStudy (anonymous):

\[\theta=\pi/6+n\pi\]

OpenStudy (anonymous):

30 degree ---> Pi/6